# Surface Reconstruction Using Rotation Systems

### v1

@@ 5.2. Synthetic datasets
In this section, we will present the reconstruction of synthetic point clouds.
@@ 5.2.1. Stanford 3D Scanning Repository
In this test, we use the vertex positions and normals from the reconstructions (Curless and Levoy, 1996) provided by Stanford Computer Graphics Laboratory.
Figure 16. Renderings of models reconstructed using our method. Top image: well-known models from the Stanford 3D Scanning Repository. Middle image: several test objects are shown. The terrain (shown after hole closing) consists of two very close surfaces, the tree exhibits significant variation point density, and the two remaining objects are high-genus shapes. Bottom: we used our method to reconstruct the DTU STL0024 model as well as the Stanford Armadillo and Bunny directly from the acquired points. For each model, we see unprocessed reconstructions on the left and models after hole closing and denoising on the right. In all images, back facing surfaces are rendered in red to show holes in the reconstruction.
Figure 17. Examples of holes in the reconstruction result on Stanford repository. The short blue lines specify the normal direction of each vertex. The left image gives the example of a hole happening on both Dragon and Happy Buddha due to bad input normal direction. Meanwhile, the right image shows the only hole on Armadillo due to an intense normal change (too sparse sampling).
Figure 16 (top) offers a qualitative demonstration of our reconstruction. All meshes are free from face intersections without considering areas with ill input. At the same time, only a few holes exist because of bad normal input (Dragon, Happy Buddha) or too sparse sampling density (Armadillo). Examples are given in Figure 17.
Table 3 illustrates practical information about reconstructing these meshes, including the number of input vertices (#.$\#.v$), running time ($t(s)$), and the number of output faces (#.$\#.f$). One can notice that the time consumption increases almost linearly with the number of vertices in the input except for Lucy. The reason why Lucy takes much longer time to finish is that it takes too much memory and hits the paging / virtual memory mechanism, which requires a lot of data transfer with disk memory. The upper bound of the input point cloud size with normal execution can be estimated. Under current hardware conditions, the maximum input size that can be processed within a linear time growth range is approximately 10M points.
@@
Model Name #.$\#.v$.$\#.f$ $t(s)$
Bunny 35,947 71,885 4
Dragon 100,250 200,143 9
Happy Buddha 144,647 288,902 14
Armadillo 172,974 345,941 16
Thai Statue 4,999,568 9,998,860 718
Lucy 14,027,812 28,055,340 13642
@@ 5.2.2. Special data
To assess the robustness of our method, we conducted an extensive experiment using a diverse set of synthetic data exhibiting special properties. Two high-genus shapes from the Thingi10k dataset (Zhou and Jacobson, 2016) were chosen to serve as a topological challenge. Additionally, a tree generated with the L-system (Prusinkiewicz et al., 1996) was employed to test the method’s robustness in handling input with significant density variations. Moreover, the inclusion of a terrain (with two sheets) and a Mobius band in the experiment aimed to evaluate the method’s performance on input with extremely close sheets. The reconstruction results on these diverse datasets are illustrated in Figure 16 (middle).
Meanwhile, comparisons are conducted with two state-of-the-art combinatorial reconstruction methods, specifically Co3Ne (Boltcheva and Lévy, 2017) (Delaunay-based) and ScaleSpace (Digne et al., 2011) (Ball-Pivoting Method), as well as the popular Screened Poisson reconstruction (Kazhdan and Hoppe, 2013), as illustrated in Fig. 18. In the case of the tree with significant density changes, Scalespace struggled to reconstruct most of the shape, while Co3Ne and Screened Poisson failed to maintain high-frequency texture on the trunk. For shapes with thin structures, such as the terrain and Mobius band, Scalespace also encountered difficulties. On the terrain dataset, our method produced a more complete mesh compared to Co3Ne, while Screened Poisson generated a watertight mesh with multiple tunnels connecting both sides. In contrast, Co3Ne achieved the best result on the Mobius band, whereas our method left a crack due to RS inconsistency around the sharp edge. Screened Poisson reconstructed a wrinkled Mobius band.
Our method, Co3Ne, and Screened Poisson performed well on the high-genus tests, while Scalespace failed the second test. Quantitative statistics are provided in Table 4. Particularly, we do not apply smoothing and any other post-processing in Co3Ne to keep the original coordinates of the vertices. Four metrics are considered including the number of unreferenced vertices ($Unref.v$), the number of intersecting faces ($Itsc.$), the number of boundary edges ($Bdry.e$), and efficiency ($t(s)$). As Screened Poisson represents a different methodological approach, meaningful insights from its statistics are challenging to extract, and thus, we omit the quantitative results for Screened Poisson.
Table 4. Quantitative result on a diverse range of data. Models in the top half are from synthetic datasets while those in the bottom half are from scanning datasets. All four indicators are better when smaller. As most meshes are not originally watertight, the number of boundary edges will barely decrease to zero. To maintain a reasonable runtime, the point cloud from DTU dataset (i.e. Pot, Animal, and Building) is reduced to approximately 500,000 points through random sampling.
Model Name #.$\#.v$ Ours Co3Ne (w/o smooth) ScaleSpace
$Unref.v$ $Itsc.$ $Bdry.e$ $t(s)$ $Unref.v$ $Itsc.$ $Bdry.e$ $t(s)$ $Unref.v$ $Itsc.$ $Bdry.e$ $t(s)$
Tree 467,740 0 64 9,620 59 674 0 21,046 0.7 168,500 284,141 4,613 627
Terrain 80,000 0 0 2,741 12 471 0 8,469 0.2 25,058 493 52,475 25
Mobius band 76,387 0 0 2,377 10 1 0 386 0.1 55,841 0 20,656 186
High-genus 1 143,662 0 0 0 21 0 0 338 0.2 50 3 0 19
High-genus 2 288,182 0 14 451 46 51 0 3720 0.5 135,659 1,491 100,850 97
Bunny 362,271 0 0 169 65 24,998 0 62,053 1.2 318 51 333 59
Armadillo 610,050 0 29 6,395 125 86,583 0 272,576 5.2 17,972 1,647 4,572 392
Pot - stl001 499,251 0 0 404 117 46,805 0 168,397 3.8 1,483 40 1,607 644
Animal - stl002 499,504 0 5 1,458 137 45,448 0 177,184 3.7 662 149 2,268 691
Building - stl024 500,000 0 11 4,594 116 25,431 0 101,693 2.4 8,418 404 2,171 831
@@
Figure 18. Visualization of the reconstruction results from four different methods on synthetic data.
@@ 5.3. Scanning datasets
For these experiments, we use the projection distance as was discussed in Section 3.5. The experiments are performed on models from the three well-known public datasets: the Stanford 3D Scanning Repository, the DTU Robot Image Data Set (randomly downsampled to 500,000 points) (Jensen et al., 2014), and the S3DIS dataset(Armeni et al., 2016) with indoor scenes. The same comparisons are performed as for the tests on synthetic data.
Figure 19 shows a comparison of our method against the three benchmark methods while Figure 27 shows the result on an acquired indoor scene. Table 4 contains the quantitative comparison. Figure 16 (bottom) shows three of the scanned models before and after denoising using bilateral normal filtering (Zheng et al., 2011).
Figure 19. This is the visualization of the reconstruction results by four different methods. It can be noticed that ours, ScaleSpace, and Screened Poisson successfully reconstruct reasonable meshes of the input point cloud, while Co3Ne fails without the smoothing operation ahead. The black area on Co3Ne is mostly another layer of surface pointing in the opposite direction as it should be. Comparing ours and ScaleSpace, our method can reconstruct a more complete mesh with a smaller number of holes while ScaleSpace is restricted by the parameter of the Ball-Pivoting algorithm.
@@ 5.4. Parameter experiment
As discussed in Section 4.5, we conducted an experiment to provide a clearer insight into the parameter $k$ by varying its values. Figure 22 showcases the reconstructed results on the raw data of the Stanford Bunny with different values of $k$. Table 5 presents corresponding statistics for running time ($t(s)$) and peak memory usage ($m(GB)$). Unsurprisingly, resource utilization increases with $k$ while we see an increase in the number of holes for small $k$.
@@ 5.6. Stress test
To investigate the robustness of our method against various common defects in input point clouds, a number of experiments were conducted. The common defects encompass noise in the point position, variations in normal direction, and misalignment.
@@ 5.6.1. Noise - point position
As we try to disentangle the influence of noise on point positions and normal directions, we compute the normal direction before the noise is added. To make the noise independent of scale, the noise vector is defined as:
@@
In addition to introducing random noise to the point positions, we further categorize the noise into two directions to assess its performance under stronger noise along both tangential and normal directions. The outcomes are depicted in Fig. 23. Remarkably, our method exhibits good robustness in the presence of noise affecting point positions.
Figure 21. Qualitative result of the ablation study. With all three tests on, the reconstructed bunny has exactly 8 holes around the bottom due to the missing data. With any of the three tests off, more small holes exist than the baseline result.
Figure 22. Qualitative result when increasing parameter $k$. The larger $k$ is, the more possible connections can be expected in the reconstructed result.
Figure 23. Noise experiment on the raw data of bunny. The first row shows the result of the experiment with noise following normal direction. The middle row depicts one in the tangential direction. While the last row is for fully random noise. The holes are rendered in red faces, it is hard to spot any holes with a glance.
@@ 5.6.2. Noise - normal direction
In this experiment, we introduce noise solely along the normal direction: this involves selecting a random direction and pivoting the normal by a constant angle $\theta$. The outcomes reveal that our method is sensitive to variations in the normal direction. When pivoting the normal away by an angle of 8 degrees, significant holes appear in the results.
Figure 24. Noise experiment on normal direction.
Figure 25. Misalignment experiment on the raw data of bunny.
@@ 5.6.3. Misalignment
Another common defect in reconstruction is misalignment. To create artificial data, we take one point cloud file from the raw data of bunny (specifically, chin.ply) as the misaligned part. The point cloud is then gradually moved away from the registered position. The length of movement is defined as:
(6) $d=A\cdot\overline{|e|}\cdot\mathbf{v}\enspace,$
where $A$ is the changing amplitude, $\overline{|e|}$ is the average edge length, and $\mathbf{v}$ is the offset direction. Fig. 25 presents the results. Our method effectively addresses misalignment when it is minimal (less than 4 times the average length). However, as misalignment increases, normal estimation is impacted, leading to the emergence of holes. When the misalignment reaches a distance 10 times the average edge length away from the original position, it is deemed to represent another layer of surface, resulting in failure.
@@
Besides displacement, other types of misalignment exist including rotational and non-rigid misalignment. Both types will create a situation where multiple layers of vertices with different normals coinhabit the same local region. If the normals are estimated from the combined point cloud, it is impossible to avoid connections between points from different layers. This leads to a situation where vertices from different layers connect with each other, often leading to holes in the reconstruction since the layers are prone to interpenetration which, of course, is a non-manifold configuration.
Examples are given through the failed reconstruction from the raw data of Stanford Happy Buddha in Fig. 26 (left). However, if the normal is accurately estimated before the alignment, no smoothing effect on normals will take place at the joint. Thus, our method will reconstruct both pieces correctly with the intersection at the joint, as the synthetic cross reconstruction shown in Fig. 26 (right). However, often the layers should be combined, and then we get a much more stable normal estimate if it is based on the combined point cloud. Hence, it is generally better to estimate the normals on the combined point cloud.
Figure 26. Left: reconstruction of the so-called Happy Buddha model from the original scan data is a fail case due to subscans which interpenetrate but still connect since points from all scans contribute to normal estimation. Right: synthetic case where a self-intersecting surface is resolved correctly by our approach by using exact normals to disambiguate.
Figure 27. The reconstructed indoor scenes from four different rooms. Most of the area, e.g. walls, chairs, table are well reconstructed.

### v2

@@ 5.1.1. Stanford 3D Scanning Repository
For the models from the Stanford Scanning Repository, we use the vertex positions and normals from the reconstructions (Curless and Levoy, 1996) provided by Stanford Computer Graphics Laboratory.
Figure 12 (top) shows the reconstructed meshes. Table 2 contains practical information about the reconstructions, including the number of input vertices ($|V|$), running time ($t(s)$), and the number of output faces ($|F|$).
@@
Model Name $|V|$ $|F|$ $t(s)$
Bunny 35,947 71,816 2
Dragon 100,240 200,268 9
Happy Buddha 144,625 289,104 15
Armadillo 172,967 345,928 15
Thai Statue 4,999,518 9,994,987 1,562
Lucy 14,025,736 28,051,203 4,301
5.1.2. Manifold ShapeNet
ShapeNet (Chang et al., 2015) is a popular dataset of 3D CAD models of objects. As many recent works use it as a main comparison dataset, we also carried out an experiment on it. However, since most meshes in this dataset are made by human designers, a large number of them are non-manifold. To perform reconstruction with reasonable input, we use the processed ShapeNet by Chu et al. (2019). We directly download the data of a processed subset of ShapeNet from this repo. Specifically, we use meshes after optimization and hidden patch removal, i.e. meshes with a suffix of _HR. There are 1,572 shapes in total, a quantitative metric of Chamfer Distance (CD) is applied to evaluate the performance. To calculate the CD score, we don’t directly use the mesh vertices. Instead, the same number of points are resampled on the reconstructed mesh. Due to resource limitations, we only compare with DSE Meshing in this experiment. To fit the best number of input points of DSE Meshing, we use Poisson disk sampling (Corsini et al., 2012) to sample 10k points with MeshLab (Cignoni et al., 2008) from the ground truth (GT) meshes. It turns out our method has better control over the topology, thus generating better reconstruction. A qualitative comparison is exhibited in Figure 13. Quantitatively, our method reaches an average CD score of 0.00836 over the 1.5k shapes, which is better than the CD score of $0.00846$ by DSE Meshing.
Figure 13. Qualitative result on Manifold ShapeNet. Non-manifold triangles are rendered red.
@@ 5.1.3. Special data
To assess the robustness of our method, we conducted experiments on a diverse collection of synthetic data sets exhibiting special properties. Two high-genus shapes from the Thingi10k dataset (Zhou and Jacobson, 2016) were chosen to serve as a topological challenge. Additionally, a tree generated using an L-system (Prusinkiewicz et al., 1996) was employed to test our method’s robustness in handling input with significant density variation. Moreover, the inclusion of a terrain (with two sheets) and a Möbius band served to evaluate the performance on inputs with extremely close sheets. The results of these experiments are shown in Figure 12 (middle).
Figure 14. Reconstructed mesh of raw bunny by DSE Meshing. Both software cannot recalculate the expected face orientation.
Comparisons are conducted with three state-of-the-art combinatorial reconstruction methods, specifically Co3Ne (Boltcheva and Lévy, 2017) (Delaunay-based), ScaleSpace (Digne et al., 2011) (Ball-Pivoting Method), and DSE meshing (Rakotosaona et al., 2021) (Machine-Learning-based), as well as the popular Screened Poisson Reconstruction (SPR) (Kazhdan and Hoppe, 2013). For the three combinatorial methods, we use the source code made available by the authors. For SPR, we use the implementation from MeshLab (Cignoni et al., 2008) with the default parameters. Also, in the interest of fairness, we generate output meshes with at least as many points as in the input data when doing SPR reconstruction. Co3Ne provides the option to apply smoothing to the input point cloud, and we include both Co3Ne with two iterations of smoothing and Co3Ne without smoothing in the comparison.
@@
Figure 15. Visualization of the reconstruction results from six different methods on special synthetic data. A detailed image enclosed by red dashed lines indicates the best reconstruction outcome.
The results of the comparisons are shown in Figure 15. Note that we render the back faces in red to highlight holes in the reconstruction. However, the output of DSE Meshing exhibits incorrectly oriented triangles, so in this case red means incorrect orientation. We tried built-in methods from Blender(Community, 2018) and MeshLab(Cignoni et al., 2008) to recalculate the orientation. However, neither application was able to correctly orient the result as illustrated in Figure 14. This indicates that DSE produces topologically flawed meshes that cannot be correctly oriented. For the results shown in Figure 15, we have arbitrarily used Blender to reorient the meshes.
Below, we provide our observations for each model regarding the relative merits of the various methods.
  • High-genus 2: On this shape, which is created by a human designer, only our method and Screened Poisson managed to accurately reconstruct the correct topology in the majority of locations (note that in places the source shape is non-manifold). The other three methods either created spurious handles or failed to reconstruct the shape.
  • Tree: This shape has multiple challenging properties. Firstly, the point density varies significantly: there are small patches with very high point density on the trunks, and the density is also much higher on the branches relative to the trunks. Secondly, the branches on the top of the tree are thin structures. In addition, there is noise along the normal direction of the surface. From the comparison, it turns out that only our method can accurately reconstruct both the textures and the general shape of the tree. Notably, DSE Meshing fails in the final triangle selection phase and does not produce an output mesh.
  • Terrain: This shape is again a thin structure with two layers pointing in opposite directions. Our method managed to reconstruct both surfaces without leaving a hole while the other methods either left holes (Co3Ne, DSE Meshing) or connected the opposite layers (Screened Poisson). Methods with smoothing (ScaleSpace, Co3Ne) will fail the reconstruction because the pre-processing mixes points from both layers. Since the neighborhood we choose to fit the tangent plane is filtered by the normal direction, our method does not suffer from the issue above.
  • High-genus 1 and Möbius band: Broadly speaking, most methods reconstruct these shapes well. That said, the methods that smooth the input points, i.e. ScaleSpace and Co3Ne, do not perform well on the Möbius band. Moreover, local topological issues arose and will be discussed below.
Quantitative statistics for the synthetic data comparisons are provided in Table 1. Three metrics are considered including vertex reference ratio ($r_{v}(\%)$), the number of boundary edges ($|E_{b}|$), and efficiency ($t(s)$). The definition of vertex reference ratio is as follows:
(3) $r_{v}=\frac{|V_{o}|}{|V_{i}|}\enspace,$
where $|V_{o}|$ is the number of vertices that are referenced in the final output mesh, and $|V_{i}|$ is the number of points as input (A referenced vertex means that it is included in at least one triangle of the mesh). This metric is used to evaluate to what extent the input information is preserved.
Because Screened Poisson Reconstruction represents a different methodological approach, we only report the number of output vertices and the time consumption in meshes generated using SPR.
Although the methods we compare against seem to perform well on Terrain, High-genus 1, and Möbius band, they do produce either incorrectly oriented parts or spurious handles in some places. We show these local issues in Figure 16.
Figure 16. Examples illustrating that Co3Ne, DSE, and SPR do fail locally on the two otherwise well-reconstructed meshes. The backfaces are here displayed in blue. In the case of Terrain, the three methods Co3Ne, DSE Meshing, and Screened Poisson connect the opposite layers. In High-genus 1, DSE Meshing generates spurious handles underneath the surface at multiple locations. A handle triangle is here shown in red.
As mentioned, volumetric methods have many advantages in terms of robustness and speed, but compared to combinatorial methods they sometimes either miss or blur subtle features. To illustrate this issue, we generated a point set with a geometric texture showing the conference logo by keeping the pixel position as the x and y coordinates while converting the pixel intensity to depth as the z coordinate. We show a comparison between our method and SPR in Figure 17. Even with the tree depth set to 15, SPR cannot reproduce the geometry accurately. While this is a synthetic example, it seems to show a scenario where combinatorial methods are preferable.
Figure 17. A synthetic example for the reconstruction of a surface with informative texture.
@@ 5.2. Scanning datasets
For these experiments, we use the projection distance as discussed in Section 3.4. The experiments are performed on models from well-known public datasets: the Stanford 3D Scanning Repository, the DTU Robot Image Data Set (randomly downsampled to around 500,000 points) (Jensen et al., 2014), Tanks and Temple(Knapitsch et al., 2017), and the S3DIS dataset(Armeni et al., 2016) with indoor scenes. The same comparisons are performed as for the synthetic data.
Figure 18. Details on noisy data reconstruction. The first row is the feet of Armadillo on the front side. The second row is a corner of the roof from Stl024.
Figure 20 shows a comparison of our method against the four benchmark methods. Specifically, DSE Meshing failed to reconstruct the Armadillo. Restricted by the size of the figure, more details are exhibited in Figure 18. Figure 19 and Figure 21 show the results on S3DIS and Tanks and Temple respectively. Table 1 contains the quantitative comparison. Figure 12 (bottom) illustrates three of the scanned models before and after denoising using bilateral normal filtering (Zheng et al., 2011b).
Figure 19. The reconstructed indoor scenes of four different rooms from S3DIS(Armeni et al., 2016). Most of the area, e.g. walls, chairs, and tables are well reconstructed.
Figure 20. This is the visualization of the reconstruction results by five different methods. It can be noticed that ours, ScaleSpace, Co3Ne w/ smoothing, and Screened Poisson successfully reconstruct reasonable meshes of the input point cloud. Comparing ours and ScaleSpace, our method can reconstruct a more complete mesh with a smaller number of holes while ScaleSpace is restricted by the parameter of the Ball-Pivoting algorithm.
Figure 21. Partial reconstruction results of dataset Tanks and Temple (Knapitsch et al., 2017). Figures with color are the rendering of the input point cloud while those in white are the reconstruction results from our method.
Figure 22. Qualitative result when increasing parameter $k$. The parameter $k$ determines the number of neighbors each vertex connects to in the initial graph $G$. The larger $k$ is, the more possible connections can be expected in the reconstructed result.
Figure 23. Qualitative result when decreasing parameter $r$. This parameter ensures that edges longer than $r$ times the average length of connections in $G$ are rejected as outliers. $k$ is set to be 50. Back faces are rendered in red.
@@ 5.3. Parameter experiment
As discussed in Section 4.5, we conducted an experiment to provide a clearer insight into the parameter $k$ by varying its values. Figure 22 showcases the reconstructed results on the raw data of the Stanford Bunny with different values of $k$. Table 3 presents corresponding statistics for running time ($t(s)$) and peak memory usage ($m(GB)$). Unsurprisingly, resource utilization increases with $k$ while we see an increase in the number of holes for small $k$. Similarly, we carried out another parameter experiment for $r$. Figure 23 demonstrates how $r$ affects the reconstruction result. Since it is used to filter out outlier connections, a small $r$ will result in a large number of edges rejected. Notably, the average edge length mentioned in Section 4.5 is related to $k$, because a larger $k$ will introduce longer connections thus increasing the average edge length.
@@ 5.5. Stress test
To investigate the robustness of our method against various common defects in input point clouds, several experiments were conducted. Common defects include noise in the point position, variations in the normal direction, and misalignment. All experiments in this section are carried out on the raw data of the Stanford Bunny, which is, of course, noisy even before the synthetic noise is added.
@@ 5.5.1. Noise - point position
In an attempt to disentangle the influence of noise on point positions and normal directions, we compute the normal direction before the noise is added. To make the noise independent of the mesh scale, the noise vector is defined as:
@@
In addition to applying random noise to all three coordinates of the point positions, we also did two experiments where noise was added separately in the tangential and normal directions. The results are shown in Figure 25 and seem to indicate that out method is quite resilient to noise applied to the positions.
Figure 25. Noise experiment on the raw data of bunny. The backface is rendered in red. The first row shows the result of the experiment with noise affecting normal direction. The middle row depicts one in the tangential direction. While the last row is for fully random noise. The holes are rendered in red faces, it is hard to spot any holes with a glance. With careful inspection, only meshes with $A=0.5$ vertical and fully random noise have small failures.
@@ 5.5.2. Noise - normal direction
In this experiment, we apply noise solely to the normal direction. This involves selecting the original normal direction as the axis of rotation and pivoting the normal by a constant angle $\theta$ around this axis. Unsurprisingly, the outcome reveals that our method is more sensitive to variations in the normal direction. When pivoting the normal away by an angle of 15 degrees, significant holes appear in the results.
This is relatively easy to explain. When a candidate edge is considered, the topology test checks if the edge belongs to a corner of the same face at either endpoint. Changing the normal confounds the rotation systems, and this means that even if a vertex belongs to the same face (as if the normals had been unperturbed) the candidate edge may not be incident on the vertex at a corner that belongs to the face.
In practice, though, this will never be an issue. The noise we applied was independent for each vertex, and this is the worst-case scenario for our reconstruction method since it relies on smoothly changing normals. Moreover, in practice, any method for normal computation would produce much smoother normal fields.
Figure 26. Noise experiment on normal direction. The normal vectors of the vertices are randomly rotated within an angle $\theta$ relative to their initial normal directions. As the noise becomes stronger, more connections are rejected, resulting in holes that are rendered in red.
Figure 27. Misalignment (displacement) experiment on the raw data of Stanford bunny. $A$ represents the magnitude of the displacement. Another layer gradually shows up as $A$ increases.
@@ 5.5.3. Misalignment
Another common defect in reconstruction is misalignment. To create artificial data, we take one point cloud file from the raw data of bunny (specifically, chin.ply) as the misaligned part. The point cloud is then gradually moved away from the registered position. The displacement is defined as:
(5) $\mathbf{d}=A\cdot\overline{|e|}\cdot\mathbf{v}\enspace,$
where $A$ is the changing amplitude, $\overline{|e|}$ is the average edge length, and $\mathbf{v}$ is the offset direction. Figure 27 presents the results. Our method effectively addresses misalignment when it is minimal (less than 4 times the average length). However, as misalignment increases, normal estimation is impacted, leading to the appearance of holes. When the misalignment reaches a distance 10 times the average edge length away from the original position, it is deemed to represent another layer of surface, resulting in failure.
@@
Besides displacement, other types of misalignment exist including rotational and non-rigid misalignment. Both types will create a situation where multiple layers of vertices with different normals coinhabit the same local region. If the normals are estimated from the combined point cloud, it is impossible to avoid connections between points from different layers. This leads to a situation where vertices from different layers connect with each other, often leading to holes in the reconstruction since the layers are prone to interpenetration which, of course, is a non-manifold configuration. Figure 28 illustrates a similar experiment on Stanford bunny. Again, chin.ply is rotated $\phi$ degrees around the z axis.
Figure 28. Misalignment (rotation) experiment on the raw data of bunny. As $\phi$ increases, two layers around the chin are reconstructed with holes and undesirable connections.
A non-synthetic example is the failed reconstruction from the raw data of Stanford Happy Buddha shown in Figure 29 (left). We note that if the normals had been correctly esitmated for each part and sufficiently different) our method would simply have meshed the different sub-scans and allowed them to intersect. A synthetic example illustrating this point is shown in Figure 29 (right). However, often the layers should be combined, and then we get a much more stable normal estimate if it is based on the combined point cloud. Hence, it is generally better to estimate the normals on the combined point cloud.
Figure 29. Left and bottom: reconstruction of the so-called Happy Buddha model from the original scan data is a fail case due to sub-scans that interpenetrate but still connect since points from all scans contribute to normal estimation. We provide the reconstruction result by Co3Ne w/ smoothing to show how challenging to reconstruct it. Right: synthetic case where a self-intersecting surface is resolved correctly by our approach by using exact normals to disambiguate.